OpenStudy (chris215):

Va is the orbital speed of a satellite at a distance x from the centre of the Earth. If the satellite is at a distance 2x from the centre of the Earth and with the speed of Vb, which of the following statements is correct?

OpenStudy (chris215):

va = vb^2?

OpenStudy (astrophysics):

 we need to use the following formula \[\huge v_{escape}^2 = \frac{ 2GM }{ r }\] so we have \[\large v_a^2 = \frac{ 2GM }{ x }\] and \[\large v_b^2 = \frac{ 2GM }{ 2x } = \frac{ GM }{ x }\]